If the earth’s orbit has a radius of 150,000,000,000 meters, what is the mass of the sun?
1
Expert's answer
2015-01-26T07:59:34-0500
m*v^2/r=G*m*M/r^2 ->M=v^2*r/G
m - mass of the Earth
M = V ^ 2 * R / G, whereM - mass of the Sun,
V - velocity of the planet 30 000 m/s,
R - radius of the orbit , (1,510^8 m)
G - gravitational constant . (6,67E- 11 m ^ 3 / (kg · s ^ 2 ))

It turns 2.02E + 30 = 2.02 × 10 ^ 30 kg .